please help me determine what M is if M=Log(8) 2+log81(27)-log(5)25+log(9)1
Answer by stanbon(48535)

M=Log(8) 2+log81(27)-log(5)25+log(9)1
-----------
M = (1/3) + (3/4) - 2 + 0
M = (13/12) -2
M= = (13-24)/12
M = -11/12
===============
Cheers,
Stan H.